A rectangle is 7 cm longer than it is wide. Find the dimensions of the rectangle if its area is 120 sq-cm
Found 2 solutions by LinnW, stanbon:
Answer by LinnW(1048)

L = length
W = width
L = W + 7
Area = (W+7)(W)
120 = W^2 + 7W
Reformatting
W^2 + 7W -120 = 0
(W + 15)(W - 8) = 0
W = -15 or W = 8
So we need to use 8 for W
Since L = W + 7
L = 15
Length is 15 and width is 8.

Answer by stanbon(75887)

A rectangle is 7 cm longer than it is wide. Find the dimensions of the rectangle if its area is 120 sq-cm
----
Let wdth = "x"
Then length = x+7
-------
Equation:
x(x+7) = 120 cm^2
--------
x^2 + 7x - 120 = 0
-------
Factor::
(x-8)(x+15) = 0
-----
Positive solution:
width = x = 8 cm
length = x+7 = 15 cm
--------------------------
Cheers,
Stan H.
=================